# Expert-Guided Prompting and Retrieval-Augmented Generation for Emergency Medical Service Question Answering

### v1

@@ Introduction
Effect of Filter. Row 6 vs. 5 measures the effect of using Filter’s predicted expertise rather than ground-truth expertise: although there is a small performance drop, ExpertRAG-Filter still surpasses the best baseline.
@@
To investigate whether our best models can be certified in NREMT exam (RQ3), we subscribed to MedicTests (MedicTests 2025), the NREMT Computer Adaptive Simulation Test. The simulation exam consists of 80-150 questions with adaptive difficulty and needs to be completed in 2.5 hours. The NREMT cognitive exam uses a scaled score to determine passing status. The passing score is 950 on a scale of 100 to 1500. We evaluated our Expert-CoT and ExpertRAG models and 0-shot baselines on the NREMT simulation exam. The expertise-augmented models use the trained filter to predict the subject area and certification. The Pass or Fail result is shown at Table 6. All models complete the exam within the allotted time with expertise-augmented models taking much longer times. Expertise-augmented models consistently lead to higher test scores and significantly improve accuracy compared to baseline LLMs. However, performance varies with model size. Specifically, 4B models fail at one or more certification levels, whereas larger 32B models pass all four. Among these, ExpertRAG‑32B with the RTF retrieval strategy achieves the highest overall score across certification levels. Importantly, although smaller models do not meet the certification threshold, their scores (940) closely approach the passing bar (950).
@@ Conclusion
This paper proposes a domain expertise-aware LLM framework for medical multiple-choice question answering that incorporates expertise to guide both LLM reasoning and RAG retrieval. We introduce EMSQA, the first large-scale labeled MCQA dataset for EMS with subject area and certification-level annotations, and curated EMS knowledge bases. We propose Expert-CoT, which guides LLM reasoning from a domain-informed perspective, and ExpertRAG, which retrieves expertise-specific knowledge for augmented generation. Experiments show that integrating expertise into LLM reasoning and RAG retrieval significantly improves performance over baselines. More importantly, our expertise-augmented models pass the NREMT simulation tests for all EMS certification levels. EMSQA provides a new benchmark for MCQA research in medical domain and our proposed expertise-augmented LLM framework can be applied to other medical MCQA datasets with similar or other expertise attributes.
@@
Although our 32B expertise-augmented model is able to pass the simulated NREMT examination, all models studied in this work are research prototypes and are not approved medical devices. They must not be used as the sole basis for diagnosis or treatment decisions for real patients. Any outputs should only be used as a reference by licensed healthcare professionals, who remain fully responsible for clinical judgment and patient care. The models may generate incorrect, incomplete, or biased recommendations and may not reflect up-to-date guidelines. Our experiments are conducted entirely in simulation, and no decisions produced by the models were used to influence real-world patient care.
@@ Acknowledgments
This work was supported by Award 70NANB21H029 from the U.S. Department of Commerce, National Institute of Standards and Technology (NIST).
Ethics Statement
Although our 32B expertise-augmented model is able to pass the simulated NREMT examination, all models studied in this work are research prototypes and are not approved medical devices. They must not be used as the sole basis for diagnosis or treatment decisions for real patients. Any outputs should only be used as a reference by licensed healthcare professionals, who remain fully responsible for clinical judgment and patient care. The models may generate incorrect, incomplete, or biased recommendations and may not reflect up-to-date guidelines. Our experiments are conducted entirely in simulation, and no decisions produced by the models were used to influence real-world patient care.
Acknowledgments
This work was supported by Award 70NANB21H029 from the U.S. Department of Commerce, National Institute of Standards and Technology (NIST).
@@ A.2.1 EMSQA Data Collection
In Tables 7, we present detailed information on 17 public and private EMSQA resources, including each source’s name and URL, number of questions, certification level and availability of explanations.
@@ A.3.2 Prompt for EMS Concept Extraction
Resource Title Level # Vocab
@@ A.3.5 Knowledge Base Collection Details
Table 10 shows the detailed information of every knowledge source we collected to construct external EMS knowledge bases. Specifically, we list the url, source type, source title, certification level, number of tokens and vocabulary of each source. In total, there are 2,545,192 tokens in our external knowledge bases, and the vocabulary size 34,110.
@@ A.4.1 NEMSIS Patient Record Preprocessing
To concatenate information for each patient records, we used the following fields in different tables, and concatenate by the shared primary key. The keys in our patient records include ”Date - Time of Symptom Onset”, ”Chief Complaints”, ”Possible Injury”, ”Cause of Injury”, ”Chief Complaint Anatomic Location”, ”Chief Complaint Organ System”, ”Gender”, ”Age”, ”Age Unit”, ”Level of Responsiveness (AVPU)”, ”Primary Symptoms”, ”Other Associated Symptoms”, ”Primary Impressions”, ”Secondary Impressions”, ”Protocol Age Category”, ”Protocols”, ”Date - Time Vital Signs Taken”, ”ECG Type”, ”SBP (Systolic Blood Pressure)”, ”Heart Rate”, ”Respiratory Rate”, ”Pulse Oximetry”, ”Blood Glucose Level”, ”End Tidal Carbon Dioxide (ETCO2)”, ”Glasgow Coma Score-Eye”, ”Glasgow Coma Score-Verbal”, ”Glasgow Coma Score-Motor”, ”Pain Scale Score”, ”Stroke Scale Score”, ”Stroke Scale Type”, ”Reperfusion Checklist”, ”Date - Time Medication Administered”, ”Medication Administered Prior to this Unit’s EMS Care”, ”Medication Given”, ”Medication Dosage”, ”Medication Dosage Units”, ”Response to Medication”, ”Role - Type of Person Administering Medication”, ”Date - Time Procedure Performed”, ”Procedure”, ”Number of Procedure Attempts”, ”Response to Procedure”, ”Role - Type of Person Performing the Procedure”, ”Alcohol - Drug Use Indicators”, ”Barriers to Patient Care”, ”Cardiac Arrest”, ”Date - Time of Cardiac Arrest”, ”First Monitored Arrest Rhythm of the Patient”, ”Cardiac Arrest Etiology”, ”Type of CPR Provided”, ”Cardiac Rhythm on Arrival at Destination”, ”Reason CPR - Resuscitation Discontinued”, ”Incident - Patient Disposition”, ”EMS Transport Method”, ”Transport Mode from Scene”, ”Initial Patient Acuity”, ”Final Patient Acuity”.
@@ A.5 Human Evaluation of EMSQA and Knowledge Base
We asked one EMT-Certified professionals rated each question–answer pair on four dimensions:
Clinical Accuracy
@@
Difficulty Level
Judge whether the question’s difficulty matches the expected level for its certification. Rating: 1 (Strongly Disagree)–5 (Strongly Agree)
Subject Area Accuracy
Confirm that the assigned medical subject area appropriately reflects the question’s content. Rating: 1 (Strongly Disagree)–5 (Strongly Agree)
Knowledge Relevance
Assess whether the assembled knowledge base sufficiently supports answering the question. Rating: 1 (Strongly Disagree)–5 (Strongly Agree)
Table 12 shows average rating for four dimensions. The results indicate that: (1) the provided answers are clinically accurate; (2) the EMT expert largely agrees with the assigned certification levels and subject areas; and (3) the curated knowledge base is helpful for answering the questions.
@@
Table 13: Retrieval Quality Across Relevance and Supportiveness
@@
A.6  Retrieval Strategy Performance
In the paper we present three different retrieval strategies (Global, Filter then Retrieve (FTR), Retrieve then Filter (RTF)), to evaluate the retrieval performance, we uniformly sampled 100 questions from different 4 certification levels (25 questions per certification level) and 10 subject areas ($\sim$ 10 questions per subject area). Then we use MedCPT to retrieve 32 documents from Knowledge Bases.
Since we don’t have the ground-truth question-document pair, we use LLM-as-a-judge (Zheng et al. 2023) to evaluate the Relevance (Es et al. 2024; Asai et al. 2023) and Supportiveness (Asai et al. 2023) of retrieved 32 documents for every question. Relevance judges if each retrieved chunk is relevant to the question. And supportiveness judges if this retrieved chunk support to answer the question.
PROMPT:
Question: {}
Figure 7: Prompt for LLM-as-a-judge
For each question $q$ in sampled questions $\mathcal{Q}$ and retrieved chunk $i$, define
Then, applied to either $\{r_{q,i}^{\mathrm{rel}}\}$ or $\{r_{q,i}^{\mathrm{sup}}\}$:
Table 13 presents the full retrieval performance of three strategies—Global (direct retrieval), Retrieve-then-Filter (RTF) and Filter-then-Retrieve (FTR)—on both “Relevance” and “Supportiveness.” In the Relevance dimension, RTF and FTR consistently outperform Global across all metrics, showing that proposed retrieval strategies retrieve relevant chunks earlier and more reliably. Likewise, in the Supportiveness dimension, both RTF and FTR achieve higher Hit@k, P@k, MRR and MAP than Global, indicating they better capture fully or partially supportive evidence. Overall, these results demonstrate that our proposed retrieval strategies more effectively extract both relevant and supportive context from the knowledge base, which in turn provides stronger grounding for the LLM’s final answer.
A7. Benchmark LLMs
A.7.1 Zero-shot Performance per Subject Area
As shown in Figure 8, we show the zero-shot LLMs’ per subject area performance on private dataset. We only run open-source models on our private dataset. The conclusion remains consistent on the private dataset: LLMs excel in pharmacology and anatomy but falter on core NREMT domains. Models answer pharmacology and anatomy items reliably, yet stumble on trauma, airway management, EMS operations, and medical&OB questions. One reason may be task complexity: the former are largely single-hop, fact-based queries, whereas the latter demand multi-hop reasoning and richer EMS context.
A.7.2 Detailed statistics of Benchmarks
Table 14 summarizes the full benchmarking results for each model and prompting strategy, reporting accuracy and F1 across all four EMS certification levels (EMR, EMT, AEMT, Paramedic) as well as overall performance. We list several key findings:
Few-shot prompting improves accuracy up to a point. We varied the number of in‑context exemplars from 0 to 64 (See Appendix Section A.7.2) and observed that incorporating a small number of examples yields substantial gain over the zero‑shot baseline. However, adding examples beyond a certain point leads to diminishing or no improvement.
Expert-CoT help guide LLM reasoning. Integrating domain expert knowledge via CoT‑Expert guides reasoning towards appropriate context and consistently boosts CoT prompting performance by up to 2.05% across models. Also, using the Filter’s predicted expertise for Expert-CoT leads to comparable performance to when using ground-truth.
@@ A.7.2 Detailed statistics of Benchmarks
Qwen3-4B RAG 78.12 79.17 75.46 76.87
+ Subject Area 81.11 81.45 79.17 80.07
+ Certification 79.16 79.21 76.38 77.25
ExpertRAG 82.24 82.26 80.51 81.16
Table 15: Ablation Study: Effect of Expertise (Certification, Subject Area) on Expert-CoT and ExpertRAG
A8. Ablation Studies on Certification and Subject Area
To assess which expertise signal contributes more to performance, we ablate two signals—Subject Area and Certification—and measure their effects on both Expert-CoT and ExpertRAG. The results are presented in Table 15.
Both expertise help, and combining them is best.. Both “Subject Area” and “Certification” help to improve performance in Expert-CoT and ExpertRAG, and combing both yields the best performance.
“Subject Area” matters more in RAG. “Subject Area” has more improvement than “Certification” in RAG. In our retrieval pipelines (RTF/FTR), “Subject Area” is used to filter irrelevant KB and PR documents, directly improving retrieved context quality. “Certification” is applied only in prompting, so it has no effect on retrieval and thus yields smaller performance gains in ExpertRAG.
A9. Error Analysis

### v2

@@ Introduction
Effect of Filter. (6) vs. (5) measures the effect of using the Filter’s predicted expertise vs. ground-truth expertise annotations. There is a small performance drop, but ExpertRAG-Filter still outperforms the best baseline.
@@
To investigate whether our best models can be certified in NREMT exam (RQ3), we subscribed to MedicTests (MedicTests 2025), the NREMT Computer Adaptive Simulation Test. The simulation exam consists of 80-150 adaptively selected questions and must be completed within 2.5 hours. The NREMT cognitive exam is scored on a 100–1500 scale, with 950 as the passing threshold. We evaluated our Expert-CoT and ExpertRAG models, along with 0-shot baselines. The expertise-augmented models used the trained Filter to predict the subject area and certification. The Pass/Fail outcomes are shown in Table 6. All models completed the exam within the allotted time, though expertise-augmented models took much longer. These models consistently achieved higher test scores and significantly improved accuracy relative to baseline LLMs. However, performance varied with model size. 4B models failed at one or more certification levels, whereas 32B models passed all four. ExpertRAG‑32B with the RTF retrieval strategy achieved the highest overall score across certifications. Notably, although the smaller LLM did not pass the test, it benefited the most from expertise augmentation, by showing the largest accuracy gains and achieving scores near or above the passing threshold.
@@ Conclusion
This paper presents a domain expertise-aware LLM framework for medical multiple-choice question answering that infers and incorporates expertise to guide LLM reasoning and RAG retrieval. We introduce EMSQA, the first large-scale labeled MCQA dataset for EMS with subject area and certification-level annotations, along with curated EMS knowledge bases. We propose Expert-CoT, which guides LLM reasoning by injecting expertise attributes into prompts, and ExpertRAG, which retrieves expertise-specific knowledge for augmented generation. Experiments show that our expertise-aware prompting and RAG strategies significantly improve performance over baselines. Importantly, the expertise-augmented LLMs pass the NREMT simulation tests across all EMS certification levels. EMSQA provides a new benchmark for MCQA research in medical domain, and our proposed expertise-aware LLM framework can be applied to other medical MCQA datasets with similar or other expertise attributes.
@@
All models studied in this work are research prototypes and not approved medical devices. They must not be used as the sole basis for diagnosis or treatment decisions. Outputs should serve only as a reference for licensed healthcare professionals, who remain fully responsible for clinical judgment and patient care. The models may generate incorrect, incomplete, or biased recommendations and may not reflect up-to-date guidelines. All experiments were conducted in simulation, with no model outputs used to influence real-world patient care. All private data were kept confidential.
@@ Acknowledgments
This work was supported by the award 70NANB21H029 from the U.S. Department of Commerce, National Institute of Standards and Technology (NIST), and a research grant from the Commonwealth Cyber Initiative (CCI).
@@ A.1 Introduction
  • A.2 EMSQA Data ........................................................................................................................................................................A.2 EMSQA Data
    • A.2.1 EMSQA Data Collection ........................................................................................................................................................................A.2.1 EMSQA Data Collection
    • A.2.2 EMSQA Data Statistics ........................................................................................................................................................................A.2.2 EMSQA Data Statistics
  • A.3 Knowledge Base Preprocessing & Statistics ........................................................................................................................................................................A.3 Knowledge Base Preprocessing & Statistics
    • A.3.1 Prompt for Organizing Chapter Text ........................................................................................................................................................................A.3.1 Prompt for Organizing Chapter Text
    • A.3.2 Prompt for EMS Concept Extraction ........................................................................................................................................................................A.3.2 Prompt for EMS Concept Extraction
    • A.3.3 UMLS Concept Normalization ........................................................................................................................................................................A.3.3 UMLS Concept Normalization
    • A.3.4 Knowledge and EMSQA Overlap Statistics ........................................................................................................................................................................A.3.4 Knowledge and EMSQA Overlap Statistics
    • A.3.5 Knowledge Base Collection Details ........................................................................................................................................................................A.3.5 Knowledge Base Collection Details
  • A.4 NEMSIS Patient Care Report Preprocessing & Statistics ........................................................................................................................................................................A.4 NEMSIS Patient Care Report Preprocessing & Statistics
    • A.4.1 NEMSIS Patient Record Preprocessing ........................................................................................................................................................................A.4.1 NEMSIS Patient Record Preprocessing
    • A.4.2 NEMSIS Patient Record Statistics ........................................................................................................................................................................A.4.2 NEMSIS Patient Record Statistics
  • A.5 Human Evaluation of EMSQA & Knowledge Base ........................................................................................................................................................................A.5 Human Evaluation of EMSQA and Knowledge Base
  • A.6 Retrieval Strategy Performance ........................................................................................................................................................................A.6  Retrieval Strategy Performance
  • A.7 Benchmark LLMs ........................................................................................................................................................................A.7 Benchmark LLMs
    • A.7.1 Zero-shot Performance per Subject Area ........................................................................................................................................................................A.7.1 Zero-shot Performance per Subject Area
    • A.7.2 Detailed statistics of Benchmarks ........................................................................................................................................................................A.7.2 Detailed statistics of Benchmarks
  • A.8 Ablation Study on Certification Level and Subject Area ........................................................................................................................................................................A.8 Ablation Study on Certification Level and Subject Area
  • A.9 Error Analysis ........................................................................................................................................................................A.9  Error Analysis
    • A.9.1 Error Types........................................................................................................................................................................A.9.1 Error Types
    • A.9.2 Error per Subject Area and Certification ........................................................................................................................................................................A.9.2  Errors per Subject Area and Certification
@@ A.2.1 EMSQA Data Collection
In Table 7, we present detailed information on 17 public and private EMSQA resources, including each source’s name and URL, number of questions, certification levels, and the availability of explanations.
@@ A.3.2 Prompt for EMS Concept Extraction
Resource Type and Title Level # Vocab
@@ A.3.5 Knowledge Base Collection Details
Table 10 shows the detailed information of every knowledge source we collected to construct external EMS knowledge bases. Specifically, we list the resource type, title, URL, certification levels, and vocabulary of each source. In total, there are 2,545,192 tokens in our external knowledge bases, and the unique vocabulary size is 34,110.
@@ A.4 NEMSIS Patient Care Report Preprocessing & Statistics
Access to NEMSIS 2021 Public-Release Research Dataset requires submitting a request on the NEMSIS website [^1] https://nemsis.org/using-ems-data/request-research-data/.
@@ A.4.1 NEMSIS Patient Record Preprocessing
To concatenate information for each patient record in the NEMSIS dataset, we used the following fields in different tables and concatenated them by the shared primary key. The keys in our patient records include ”Date - Time of Symptom Onset”, ”Chief Complaints”, ”Possible Injury”, ”Cause of Injury”, ”Chief Complaint Anatomic Location”, ”Chief Complaint Organ System”, ”Gender”, ”Age”, ”Age Unit”, ”Level of Responsiveness (AVPU)”, ”Primary Symptoms”, ”Other Associated Symptoms”, ”Primary Impressions”, ”Secondary Impressions”, ”Protocol Age Category”, ”Protocols”, ”Date - Time Vital Signs Taken”, ”ECG Type”, ”SBP (Systolic Blood Pressure)”, ”Heart Rate”, ”Respiratory Rate”, ”Pulse Oximetry”, ”Blood Glucose Level”, ”End Tidal Carbon Dioxide (ETCO2)”, ”Glasgow Coma Score-Eye”, ”Glasgow Coma Score-Verbal”, ”Glasgow Coma Score-Motor”, ”Pain Scale Score”, ”Stroke Scale Score”, ”Stroke Scale Type”, ”Reperfusion Checklist”, ”Date - Time Medication Administered”, ”Medication Administered Prior to this Unit’s EMS Care”, ”Medication Given”, ”Medication Dosage”, ”Medication Dosage Units”, ”Response to Medication”, ”Role - Type of Person Administering Medication”, ”Date - Time Procedure Performed”, ”Procedure”, ”Number of Procedure Attempts”, ”Response to Procedure”, ”Role - Type of Person Performing the Procedure”, ”Alcohol - Drug Use Indicators”, ”Barriers to Patient Care”, ”Cardiac Arrest”, ”Date - Time of Cardiac Arrest”, ”First Monitored Arrest Rhythm of the Patient”, ”Cardiac Arrest Etiology”, ”Type of CPR Provided”, ”Cardiac Rhythm on Arrival at Destination”, ”Reason CPR - Resuscitation Discontinued”, ”Incident - Patient Disposition”, ”EMS Transport Method”, ”Transport Mode from Scene”, ”Initial Patient Acuity”, ”Final Patient Acuity”.
@@ A.5 Human Evaluation of EMSQA and Knowledge Base
We asked one EMT-certified professional to rate each question–answer pair on four dimensions:
Clinical Accuracy:
@@
Difficulty Level:
Judge whether the question’s difficulty matches the expected level for its certification. Rating: 1 (Strongly Disagree) to 5 (Strongly Agree)
Subject Area Accuracy:
Confirm that the assigned medical subject area appropriately reflects the question’s content. Rating: 1 (Strongly Disagree) to 5 (Strongly Agree)
Knowledge Relevance:
Assess whether the assembled knowledge base sufficiently supports answering the question. Rating: 1 (Strongly Disagree) to 5 (Strongly Agree)
Table 12 shows average rating for four dimensions. The results indicate that: (1) provided answers are clinically accurate; (2) the EMT expert largely agrees with the assigned certification levels and subject areas; and (3) the curated knowledge base is helpful for answering the questions.
@@
A.6  Retrieval Strategy Performance
In the paper we present three different retrieval strategies (Global, Filter then Retrieve (FTR), Retrieve then Filter (RTF)). To evaluate the retrieval performance, we uniformly sampled 100 questions from the 4 different certification levels (25 questions per certification level) and 10 subject areas ($\sim$ 10 questions per subject area). For each question, we use MedCPT to retrieve 32 documents from Knowledge Bases.
Since we do not have the ground-truth question-document pairs, we use LLM-as-a-judge (Zheng et al. 2023) to evaluate the “Relevance” (Es et al. 2024; Asai et al. 2023) and “Supportiveness” (Asai et al. 2023) of the 32 retrieved documents for every question. Relevance judges if each retrieved chunk is pertinent to the question, while supportiveness judges if this retrieved chunk helps answer the question.
PROMPT:
Question: {}
Figure 7: Prompt for LLM-as-a-judge
For each question $q$ in sampled questions $\mathcal{Q}$ and retrieved chunk $i$, we define:
Then, for either $\{r_{q,i}^{\mathrm{rel}}\}$ or $\{r_{q,i}^{\mathrm{sup}}\}$:
@@ A.6  Retrieval Strategy Performance
Table 13: Retrieval Relevance and Supportiveness
Table 13 presents the performance of three retrieval strategies,Global (direct retrieval), Retrieve-then-Filter (RTF) and Filter-then-Retrieve (FTR), in terms of both “Relevance” and “Supportiveness”. In terms of “Relevance”, RTF and FTR show slightly better performance than Global across all metrics, indicating that the proposed strategies retrieve relevant chunks earlier and more reliably. In terms of “Supportiveness”, both RTF and FTR achieve higher Hit@k and P@k and slightly better MRR and MAP than Global, suggesting they more effectively capture fully or partially supportive evidence. Overall, although based on a small study, these results show that the proposed retrieval strategies improve the extraction of both relevant and supportive context from the knowledge base, providing stronger grounding for the LLM’s final answer. A more comprehensive evaluation of retrieval strategies is beyond the scope of this paper and left for future work.
A.7 Benchmark LLMs
A.7.1 Zero-shot Performance per Subject Area
As shown in Figure 8, we show the zero-shot LLMs’ per subject area performance on private dataset. We only run open-source models on our private dataset. The conclusion remains consistent on the private dataset: LLMs excel in pharmacology and anatomy but falter on core NREMT domains. Models answer pharmacology and anatomy items reliably, yet stumble on “trauma”, “airway management”, “EMS operations”, and “medical&OB” questions. One reason may be task complexity: the former are largely single-hop, fact-based queries, whereas the latter demand multi-hop reasoning and richer EMS context.
A.7.2 Detailed statistics of Benchmarks
Table 14 summarizes the full benchmarking results for each model and prompting strategy, reporting accuracy and F1 across all four EMS certification levels (EMR, EMT, AEMT, Paramedic) as well as overall performance. Below are several key findings:
Few-shot prompting improves accuracy up to a point. We varied the number of in‑context exemplars from 0 to 64 and observed that incorporating a small number of examples yields substantial gain over the zero‑shot baseline. However, adding examples beyond a certain point leads to diminishing or no improvement.
Expert-CoT helps guide LLM reasoning. Integrating domain expert knowledge via CoT‑Expert guides reasoning towards appropriate context and consistently boosts CoT prompting performance by up to 2.05% across models. Also, using the Filter’s predicted expertise for Expert-CoT leads to comparable performance to when using ground-truth.
@@ A.7.2 Detailed statistics of Benchmarks
Qwen3-4B CoT 72.35 73.09 70.58 72.02
+ Subject Area 76.32 77.04 73.82 74.53
+ Certification 74.53 75.37 73.29 74.31
Expert-CoT 77.26 78.38 74.32 75.77
Table 15: Ablation Study: Effect of Expertise (Certification, Subject Area) on Expert-CoT
A.8 Ablation Study on Certification Level and Subject Area
To determine which expertise attribute contributes the most to performance, we ablate the two key attributes, “Subject Area” and “Certification”, and measure their individual effects on Expert-CoT. We cannot do the ablation study on Expert-RAG since it only uses the subject area for retrieval. The results are presented in Table 15.
Expertise attributes are most effective when used together. Individually, “Subject Area” and “Certification” improve performance in Expert-CoT between 0.74-4.91% in Acc, but combining both attributes yields the best improvements of up to 5.29% in the 4B model. Besides, “Subject Area” helps slightly more than “Certification” in performance improvement.
Small language models benefit more from expertise guidance. Injecting expertise labels yields only marginal improvements for the 32B LLM model, but leads to significant gains for the 4B model. This disparity suggests that larger models have likely internalized much of the relevant expertise during pre-training, whereas smaller models rely more heavily on explicit expertise guidance to steer their limited capacity toward the correct reasoning trajectory.
A.9  Error Analysis
A.9.1 Error Types
We conducted an error analysis on our best 4B model (ExpertRAG-GT + Expert-CoT), which achieved 82.24% accuracy on EMSQA public split. From the 661 errors on the test set, we randomly selected 50 samples and manually examined each question, its retrieved knowledge, and the model’s reasoning process. As shown in Figure 9, we attribute the errors to five main categories:
1) Reasoning Error (42.4%): Retrieved documents were relevant and were used to answer the question, but the model’s reasoning was incorrect.
2) Retrieval Error(20.3%): No relevant knowledge was retrieved, causing the model to rely on internal “model knowledge”, which led to incorrect answer.
3) Knowledge Use Failure (20.3%): Part of the retrieved knowledge was relevant to answer the question but the model used other irrelevant knowledge in its reasoning.
4) Question Misinterpretation (15.3%): The model misunderstood the question. For example, for the question what is the age cutoff for using an AED on a patient? with choices: a. 18 year old; b. 5 year old; c. 25 year old; d. 1 year old, the question refers to minimum patient age (Correct answer: d. 1 year old) for routine AED use. However the model interpreted it as the minimum age of the responder allowed to operate an AED and incorrectly selected (a. 18 year old).
5) Model Policy Constraint (1.7%): The model determined that none of the answer choices were valid and abstained from responding.